Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

probability question

Expert replies
by robk100 » Fri Dec 14, 2007 3:31 pm
Hi for some reason i'm having trouble comprehending this.

I want to roll one dice. The first roll must be an even number. The second roll must be the number 4. I know the chances of doing this is 1/12, but what's the chances if the order doesn't matter. Ie. whether i roll a 4 or even number first doesn't matter as long as the second roll completes it. Please explain the answer and why.

Thanks in advance
Join the discussion
Source: — Problem Solving |

Re: probability question

by gabriel » Sun Dec 16, 2007 12:23 pm
robk100 wrote:Hi for some reason i'm having trouble comprehending this.

I want to roll one dice. The first roll must be an even number. The second roll must be the number 4. I know the chances of doing this is 1/12, but what's the chances if the order doesn't matter. Ie. whether i roll a 4 or even number first doesn't matter as long as the second roll completes it. Please explain the answer and why.

Thanks in advance
When the order doesnt mater there are two ways this event could happen, first you get a 4 and then you get a even number OR first u get a even number and then you get a 4, the probability of each of them happening is 1/12 so the total probability = 1/12+1/12 = 1/6 ..

Regards.
Join the discussion

thank you

by robk100 » Sun Dec 16, 2007 4:00 pm
Thank you for your reply I understand it now....
Join the discussion